OpenStudy (anonymous):

For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 2x + y = 9 −x − y = −8 −8 + y −8 − y 8 + y 8 − y

OpenStudy (anonymous):



OpenStudy (anonymous):

-x-y=-8 Multiply both sides by -1 then, x+y=8 Subtract y from both sides then, x+y-y=8-y x=8-y

OpenStudy (anonymous):

got it?